OpenStudy (anonymous):

Find the number of 5 card hands that can be drawn from a deck of 52 cards

OpenStudy (zarkon):

a Combination

jimthompson5910 (jim_thompson5910):

There are 52 cards for the 1st draw 51 cards for the 2nd draw 50 cards for the 3rd draw 49 cards for the 4th draw 48 cards for the 5th draw So there are 52*51*50*49*48 = ??? ways total to draw 5 cards IF ORDER MATTERS Since order does NOT matter, this means that we must divide the result by 5! = 5*4*3*2*1 = 120 to remove any duplicates.

jimthompson5910 (jim_thompson5910):

I should say "52 cards for the 1st draw" really means "52 cards to choose from for the 1st draw" and this rewording applies to each line from 52 on down to 48

OpenStudy (anonymous):

so I would multiply 52 5 times?

jimthompson5910 (jim_thompson5910):

no, not quite

jimthompson5910 (jim_thompson5910):

that would only be the case if you replaced each card you drew

OpenStudy (anonymous):

so i would divide

jimthompson5910 (jim_thompson5910):

you would first compute 52*51*50*49*48

jimthompson5910 (jim_thompson5910):

then you would divide by 120

OpenStudy (anonymous):

so 2598960 would be the answwe

OpenStudy (anonymous):

*answer

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thank you:)

jimthompson5910 (jim_thompson5910):

np